KR’96

Fifth International Conference on Principles of Knowledge Representation and Reasoning

KR’96

Cambridge, Massachusetts, U.S.A.  World Wide Web: http://www.kr.org/kr/
November 5–8, 1996  Autoresponder: kr96-info@kr.org

Explicit representations of knowledge manipulated by inference algorithms provide an important foundation for much work in Artificial Intelligence, from natural language dialogue systems to expert systems. We intend KR’96 to be a place for the exchange of news, issues, and results among the community of researchers in the principles and practices of knowledge representation and reasoning (KR&R) systems.

We encourage papers that present substantial new results in the principles of KR&R systems while clearly showing the applicability of those results to implemented or implementable AI systems. We also encourage “reports from the field” of applications, experiments, developments, and tests. The following topics are meant to be suggestive of the scope of the conference.

Representational Formalisms Representations of Belief,

Intention Time, Space, Action, Events

Nonmonotonic Logics Description Logics

Reasoning Techniques Deduction Induction Abduction Reasoning under Uncertainty Parallel and Distributed Implementations Efficiency Measures and Complexity Implemented KR&R Systems

Reports, Updates

Comparisons, Evaluations

Significant Applications Planning Robotics Diagnosis Natural Language Multi-Agent Environments Knowledge Bases

Implications for/of Machine Learning Decision Theory Databases Software Engineering

Full details available via the web page and autoresponder email address given above.

SCHEDULE

May 6, 1996  Extended abstracts due
July 1, 1996  Results to authors
August 15, 1996  Final papers due
November 5-8, 1996 KR’96

CONFERENCE COMMITTEE

Conference Chair  Jon Doyle  doyle@mit.edu
Program Co-Chairs  Luigia Carlucci Aiello aiello@dis.uniroma1.it

                                    Stuart C. Shapiro  shapiro@cs.buffalo.edu

Inter-Conference Cooperation Chair Ronald P. Loui  loui@cs.wustl.edu
Publicity Chair  Werner Horn  werner@ai.univie.ac.at

KR’96 will be held immediately preceding the AAAI Fall Symposia Series held Nov. 9–11 (see http://www.aaai.org/), and immediately after several independent workshops tentatively held Nov. 2–4, all also in Cambridge: Description Logic ’96 (contact dl96@dl.kr.org), and Relevance in Knowledge Representation and Reasoning (contact levy@research.att.com

or greiner@scr.siemens.com).